# Kramer’s Escape Rate and Phase Transition Dynamics in AdS Black Holes with Dark Structures

### v1

@@ Abstract
Traditional static methods in phase transition studies offer insights into black hole thermodynamics but miss dynamic aspects. Kramers’ escape rate, central to our research, has a somewhat dynamic approach to phase transition. We explore free energy landscapes for black holes influenced by ’dark’ and ’dark +strings’ entities, assessing how additional parameters affect escape and transition rates during phase transitions from the small black holes to the large black holes. In our dual-perspective analysis, we first consider the escape rate as a function of the black hole radius, demonstrating a natural increase in the escape rate up to a Maximum point, followed by a decrease due to a reaction mechanism. From this viewpoint, the escape rate from smaller to larger black holes will consistently be greater than the reverse. From the second perspective, we consider the escape rate as a function of temperature (pressure). In both models under study, initially, the behavior of the probability diagrams is akin to the radius-based behavior. However, with an increase in temperature (and corresponding pressure), we reach a specific and distinct point, where the systems will move towards either smaller or larger black holes with the same probability. Beyond this temperature (pressure), the behavior of the models diverges. In the nonlinear model, an inversion occurs where the escape rate from larger black holes to smaller ones becomes predominant. This inversion area seems to be significantly reduced compared to the charged black hole model, although it is still discernible in the temperature-based escape rate diagram. In other words, within the temperature range studied, where the likelihood of radiative conditions and movement from a large black hole to a small black hole is less, It seems that the addition of dark parameters minimizes the occurrence of this transition, indicating the subtle influence of dark structures on the phase transition dynamics between black holes. In the string model, although the trend of increasing pressure in the transition from a large black hole to a smaller one is greater than the reverse, it seems that the transition process from a small black hole to a larger one, with decreasing pressure, completely prevents the ’low probable’ transition from a large black hole to a smaller one. This is an interesting and distinctive observation.
@@ 2.1 Kramer’s escape rate
Let us consider the the Fokker-Planck equation [67, 70, 71],
@@ 2.2.1 Gibbs Free Energy Landscape
During the intricate process of phase transition, particularly of the first order, an assortment of black holes of varying sizes-small, medium, and large-along with transitional states, frequently emerges. Consequently, it is feasible to conceptualize a collection of these black hole space-times, each with an arbitrary horizon radius, at a specified temperature ”T”. For a better understanding, it must be stated that one of the most important characteristics of a black hole is its temperature, or more specifically, the Hawking temperature. The mere existence of temperature, regardless of how it is defined, implies the possibility of thermal fluctuations within the system. Consider this as a background. Now, the black hole under study has been excited for some environmental reason and begins a phase transition process. At both the initial and ends of this phase transition, there are two locally stable black holes, one of which is globally more stable than the other, setting the stage for this transition. However, an important point regarding these two states is that both must necessarily be solutions to Einstein’s field equations. The golden key to defining a set of black holes lies here. Although the system must end up in a state that is necessarily a solution to Einstein’s equations at both ends, there is no need to strictly follow the static field equations during the transition from the initial phase to the final phase, since the system is in an unstable transitional phase. Therefore, considering a predefined energy range and with the condition that the chosen paths do not destroy the overall stability of the system, the system can be allowed to traverse any desired path from the origin to the end. This freedom of choice and multiple paths could serve as a gateway for the introduction of probability and statistical methods. Hence, by selecting an arbitrary indistinctive set within this framework, which acts as a canonical ensemble for us, a statistical framework for analysis can be presented. It is noteworthy that the temperature ”T” of the ensemble may differ from the Hawking temperature associated with black holes Notably, the temperature ”T” of the ensemble may diverge from the Hawking temperature associated with black holes. Each of the space-time configurations within this ensemble is characterized by a distinct Gibbs free energy. Hence, we refer to this construct as the generalized Gibbs result or, more descriptively, the Gibbs free energy landscape.
@@ 2.2.2 Landau free energy
The Landau free energy constitutes another showy path for scrutinizing the phase transition dynamics within the system’s free energy landscape. While using this method assumes heightened significance in the context of second-order phase transitions- wherein the bifurcation of the Landau functional’s local minimum epitomizes the thermodynamic system’s second-order phase transition- it is also pertinent to first-order phase transitions. Specifically, the transmutation of the global minimum is emblematic of a first-order phase transition, underscoring the multifaceted nature of the Landau-free energy in characterizing phase behavior. This function can be written as follows, [57]
@@
Given that ”L” is an energy function, it inherently embodies the fundamental characteristic of energy minimization. This implies that the system invariably seeks a trajectory that reduces the energy expenditure required for transformation. Therefore, among the triad of system parameters $(X,P,T)$ that influence the ”F” function, only the configurations that minimize ”L” will accurately reflect the system’s state.
@@ 3.0.1 Behavioral comparison of landscape free energy
Now according to equations (2), (3), (4) and (7) for Gibbs and Landau free energy function and also the thermal potential of this model, we will have,
@@
Before any calculations, it is better first compare the various forms of free energy definitions, namely $G_{L}$, U, and L. Theoretically, from the canonical perspective the number of particles and mass should remain constant, consequently only the entropy-temperature and pressure-volume terms will play a role in the first law of thermodynamics. Additionally, given that in Landau’s free energy, $X$ can be considered equivalent to volume, it appears that ultimately, all three introduced forms for the free energy landscape must coincide. In Fig (2a) and (2b), for the model under study, we have plotted these three definitions once according to their general structure and once based on their dimensionless state. Practically, in terms of the shape structure, all three forms have the same trend, but there are also some visible differences, some of which may be attributed to computational precision. Moreover, since Landau’s free energy has dimensions of energy, we used the $G_{c}$ to render it dimensionless, which seems to have increased the degree of displacement, although the shape has been preserved. An intriguing point that comes to mind before concluding this discussion is that in statistical ensembles, such as the grand canonical ensemble where terms related to the chemical potential play a more prominent role, the complexities and structural differences of these three definitions may become more apparent compared to this scenario. Concerning the above discussion and as seen in Fig (2), all the shapes of the free energy landscape (with a little tolerance for displacement) will practically have the same shape structure in the canonical view. Accordingly, the choice and use of each do not seem to have an advantage over the others. For this reason, we will henceforth utilize them under the designation of the free energy landscape, without loss of generality.
@@ 3.0.2 Kramer’s escape rate
An important point to focus on initially is the difference in concept between this section and the previous one. Contrary to the previous section, here we do not intend to study the influence of temperature on free energy and the process of transition from a small to a large black hole during a first-order phase transition. Instead, we aim to examine, in light of the concepts previously discussed for Kramer’s escape rate -which represents the rate of particle escape under thermal fluctuations from a known local minimum in the potential- whether conditions might arise that could reverse the transition process from a small to a large black hole? Typically, we expect a transition from an unstable local minimum to a stable global minimum. However, the question arises whether there is a range in which, due to environmental conditions such as changes in temperature or pressure, the probability of a reverse process, namely the transition from a large to a small black hole, could precede the primary and natural process, causing systems to move in reverse from the large to the small black hole. To better understand, Fig (5a) illustrates a specific instance of a free energy landscape at a defined pressure of $p=0.875692$. For this particular scenario, we examine two states:
1. The investigation of leakage or the probability of transition from the unstable local minimum alpha to the stable global minimum gamma, which is conventionally expected to occur with an increase in temperature.
2. The examination of the likelihood that, under certain environmental conditions such as increased temperature or pressure, the reverse process, namely the transition from gamma to alpha, could become the dominant process, even if only in a small region.
Figure 5: The graph of $G_{L}$ against x for free parameters in fig (5a) and T versus x in fig (5b) which are determined as the locations of ($\alpha$) intersection (red and orange), ($\beta$) intersection (blue and orange) and ($\gamma$) intersection (green and orange)
Now, using equation (1), we calculate the escape rate of each of the possible transitions, i.e. ($\alpha\rightarrow\beta$) and ($\gamma\rightarrow\beta$), the result of which can be seen in figure (6).
Figure 6: (6a): $r_{k}$ with respect to $x$,  (6b): $r_{k}$ with respect to $t$ and (6c): Enlarged detail, bottom part of diagram (6b)
In Figure 6a, we have plotted the escape rate as a function of distance. For a better interpretation of this graph, it is advisable to read it concurrently with the energy versus distance graph fig (5a). Initially, as energy increases from $\alpha$ towards $\beta$, we should witness an increase in the escape growth, which, as confirmed by both figures, should peak at a distance of 1.014. Then, with a gradual decrease in energy from $\beta$ towards $\gamma$, we should observe a decreasing trend in the probability. Considering the relatively significant height difference between $\gamma\rightarrow\beta$ and $\alpha\rightarrow\beta$, we expect that throughout the entire range studied, the dominant process should always be the transition from a small to a large black hole, which the graph precisely confirms our classical understanding.
In Figure 6b, we have depicted the escape rate as a function of temperature t. At first glance, the process appears similar to the previous case; that is, initially, with an increase in temperature and considering the small radius, we should observe an increase in the escape rate until it reaches a peak, and then, despite the temperature increase, due to the radius enlargement and energy decrease, we should witness a decreasing trend in probability, which the shape of the graph confirms. However, an analysis of the end section of this graph, which we have illustrated in Figure 6c, reveals that contrary to the previous case, it represents a completely different situation within a small range.
At ( t = 0.8976 ), the two escape rate curves intersect, which, makes their occurrence probabilities equivalent. This means that at this temperature, black holes have the discretion to choose whether to follow the classic process from a small to a large black hole or to undertake a reverse process. Subsequently, it is the reverse process that becomes dominant, meaning that, although this range may seem exceedingly small, the system nonetheless exhibits a propensity within this limited interval to transition from a large black hole to a small black hole, a tendency that may seem contrary to classical expectations.
@@
where q is the charge of the black hole, ”a” is related to the density of clouds of strings and comes from the energy-momentum tensor of the clouds of strings and g, which is displayed with g instead of $\alpha$ to avoid mistakes with clouds of strings parameter, is a positive Gauss-Bonnet coupling constant.
@@ 4.0.2 Kramer’s escape rate
In this black hole, unlike the previous case where we examined a specific state of the potential of the system at a certain pressure in terms of temperature changes, we are looking to investigate the changes in the rate of escape, at a constant temperature for a certain state of the system, in terms of pressure. For this purpose, we consider the temperature t = 0.937, which is equivalent to the energy state of the system in Fig (7a).
Figure 7: The graph of $G_{L}$ against x for free parameters in fig (7a) and p versus x in fig (7b) which are determined as the locations of ($\alpha$) intersection (red and orange), ($\beta$) intersection (blue and orange) and ($\gamma$) intersection (green and orange)
Now, using equation (1), we calculate the escape rate of each of the possible transitions, i.e. ($\alpha\rightarrow\beta$) and ($\gamma\rightarrow\beta$), the result of which can be seen in fig (8).
Figure 8: (8a): $r_{k}$ with respect to $x$,  (8b): $r_{k}$ with respect to $p$ and (8c): Enlarged detail, bottom part of diagram (8b)
As it is evident from Fig (8c), at (p = 0.8384), again the two escape rate curves intersect, which makes their occurrence probabilities equivalent. It appears that after this convergence point, a process entirely different from the nonlinear or charged model takes shape. This means that quantitatively, the pressure associated with the transition from a large black hole to a small one increases more than in the reverse case. However, this increase in pressure for the transition from a large to a small black hole does not actually cause the corresponding probability to increase as well, leading not to the dominance of the large to small black hole trend. As can be seen from the figure, if we draw an isobar, we find that $r_{k1}$ remains larger than $r_{k2}$. This may indicate a greater harmony and alignment of this model with the known natural laws of physics
@@ 5 Discussion and comparison
As observed in Figs (5a) and (7a), our black hole initially resides in a stable local minimum, alpha, which globally is considered an unstable minimum. Consequently, as one would naturally expect, with the slightest thermal perturbation, and corresponding pressure, the system increasingly tends to transition towards the global minimum,$\gamma$, signifying an increase in radius. As the radius increases and the system begins to move towards a larger black hole, according to the reaction principle, it is natural that processes that prevent excessive radial growth and uncontrolled instability also form. This leads to the fact that the escape rate after reaching a peak, a decreasing trend will ensue, and the system will potentially stabilize at the global minimum (Fig (6c)).
However, in (NLM-C-Q-PFD) modle, when examining the escape rate diagram as a function of temperature (or pressure), we find that when the temperature (and corresponding pressure) exceeds a certain limit, a small range emerges where a reverse process forms and prevails Fig (6c). In other words, With $r_{k2}$ prevailing over $r_{k1}$, a situation is created that allows black holes to move from $\gamma$ to $\beta$. It seems that the transition from $\gamma$ to $\beta$ is less consistent with physical conditions. This is because, within the same energy structure and without topological changes in the energy function (that is, maintaining the gamma state as the global minimum and alpha as the local minimum), the system prefers to reach an $\alpha$ state with a smaller radius and consequently reduced entropy, which is at a higher energetic level. Therefore, from our perspective, this situation can be termed ’low probable.’ In our opinion, based on the results, it appears that the larger this region is, meaning the reverse process begins at lower temperatures or pressures, it may indicate an incomplete structural evolution of the black hole model to fully conform with the laws of nature. For example, in the Reissner-Nordstrm black hole [48], as a basic model that includes only electric charge, this region is significantly larger and practically quite evident. As we know, due to the choice of mass and charge size, this model has repeatedly shown non-conformity with standard physical conditions and cosmological conjectures [61,62]. In other words, from comparing the results of the Reissner-Nordstrm model with the (NLM-C-Q-PFD) black hole, it seems that the addition of parameters influenced by dark structure (energy and matter) has led to the black hole behavior approaching a more natural and orderly structure. The situation regarding the (EGB-YM-CS) structure is completely different from the previous state. It appears that after the convergence point of the two probability diagrams, which gives black holes an equal choice to move between two minima, the transition from $\gamma$ to $\beta$, in terms of increasing pressure changes, overtakes the transition from $\alpha$ to $\beta$. However, this increase in pressure, according to diagram (8c), is not accompanied by an increased probability of transition; that is, if we draw an isobar in any region after the convergence, we will see that the transition probability $r_{k1}$ remains greater than $r_{k2}$ It seems that the transition from $\alpha$ to $\beta$ after the convergence point, with decreasing pressure, prevents the ’low probable’ process and displays a complete harmony with nature.

### v2

@@ Abstract
Traditional static methods in phase transition studies, such as the swallowtail bifurcation diagram, provide good insights into the thermodynamics of black holes. However, they practically lose sight of the dynamic aspects and temporal sequence of events. The Kramers escape rate, central to our research, offers a somewhat dynamic approach to phase transition. We examine the free energy landscapes for black holes under the influence of ’dark’ and ’stringy+dark’ structures, assessing how additional parameters affect the escape rates and dynamics of the transition during the first-order phase transition from small to large black holes. In our analysis, we consider the escape rate as a function of the black hole radius and study its variations. We will observe that, on one hand, the escape rate well represents our assumption based on the movement from zero, increasing to a maximum point, and then decreasing back to zero as reactive structures become active during the phase transition interval. However, the critical point in this method is the encounter with a specific and distinct point. This is where the diagram of the direct process (escape rate from small to large black holes) intersects with the reverse process (large to small black holes), becoming equally probable (contact point). The point, which seems improbable at the onset of the phase transition or very negligible, gains more significance as the process progresses. This increase indicates the dominance of a region where the escape rate from larger black holes to smaller ones prevails. The predominance of the reverse process, which increases as we approach the end of the process and is necessarily accompanied by a variation in radius, may be considered as a natural reaction of the black hole against the ’phase change’ action. A reaction which attempting to prevent any uncontrolled radial growth that could jeopardize the stability of the black hole.
@@ 2.1 Kramer’s escape rate
Let us consider the Fokker-Planck equation [67, 70, 71],
@@ 2.2.1 Gibbs Free Energy Landscape
During the intricate process of phase transition, particularly of the first order, an assortment of black holes of varying sizes-small, medium, and large-along with transitional states, frequently emerges. Consequently, it is feasible to conceptualize a collection of these black hole space-times, each with an arbitrary horizon radius, at a specified temperature ”T”. For a better understanding, it must be stated that one of the most important characteristics of a black hole is its temperature, or more specifically, the Hawking temperature. The mere existence of temperature, regardless of how it is defined, implies the possibility of thermal fluctuations within the system. Consider this as a background. Now, the black hole under study has been excited for some environmental reason and begins a phase transition process. At both the initial and ends of this phase transition, there are two locally stable black holes, one of which is globally more stable than the other, setting the stage for this transition. However, an important point regarding these two states is that both must necessarily be solutions to Einstein’s field equations. The golden key to defining a set of black holes lies here. Although the system must end up in a state that is necessarily a solution to Einstein’s equations at both ends, there is no need to strictly follow the static field equations during the transition from the initial phase to the final phase, since the system is in an unstable transitional phase. Therefore, considering a predefined energy range and with the condition that the chosen paths do not destroy the overall stability of the system, the system can be allowed to traverse any desired path from the origin to the end. This freedom of choice and multiple paths could serve as a gateway for the introduction of probability and statistical methods. Hence, by selecting an arbitrary indistinctive set within this framework, which acts as a canonical ensemble for us, a statistical framework for analysis can be presented. Notably, the temperature ”T” of the ensemble may diverge from the Hawking temperature associated with black holes. Each of the space-time configurations within this ensemble is characterized by a distinct Gibbs free energy. Hence, we refer to this construct as the generalized Gibbs result or, more descriptively, the Gibbs free energy landscape.
@@ 2.2.2 Landau free energy
The Landau free energy constitutes another showy path for scrutinizing the phase transition dynamics within the system’s free energy landscape. While using this method assumes heightened significance in the context of second-order phase transitions- wherein the bifurcation of the Landau functional’s local minimum epitomizes the thermodynamic system’s second-order phase transition- it is also pertinent to first-order phase transitions. Specifically, the transmutation of the global minimum is emblematic of a first-order phase transition, underscoring the multifaceted nature of the Landau-free energy in characterizing phase behavior. This function can be written as follows [57],
@@
Given that ”L” is an energy function, it inherently embodies the fundamental characteristic of energy minimization. This implies that the system invariably seeks a trajectory that reduces the energy expenditure required for transformation. Therefore, among the triad of system parameters $(X,P,T)$ that influence the ”F” function, only the configurations that minimize ”L” will accurately reflect the system’s state,
@@ 3.0.1 Behavioral comparison of landscape free energy
Now according to equations (2), (3), (4) and (7) for Gibbs and Landau free energy function and also the thermal potential of this model, we will have [^5] Note that in the following relationships, due to the number of parameters, the beginning of the equations is shown with a bold symbol only for better clarity.,
@@
Before any calculations, it is better first compare the various forms of free energy definitions, namely $G_{L}$, U, and L. Theoretically, from the canonical perspective the number of particles and mass should remain constant, consequently only the entropy-temperature and pressure-volume terms will play a role in the first law of thermodynamics. Additionally, given that in Landau’s free energy, $X$ can be considered equivalent to volume, it appears that ultimately, all three introduced forms for the free energy landscape must coincide. In Figs (2a) and (2b), for the model under study, we have plotted these three definitions once according to their general structure and once based on their dimensionless state. Practically, in terms of the shape structure, all three forms have the same trend, but there are also some visible differences, some of which may be attributed to computational precision. Moreover, since Landau’s free energy has dimensions of energy, we used the $G_{c}$ to render it dimensionless, which seems to have increased the degree of displacement, although the shape has been preserved. An intriguing point that comes to mind before concluding this discussion is that in statistical ensembles, such as the grand canonical ensemble where terms related to the chemical potential play a more prominent role, the complexities and structural differences of these three definitions may become more apparent compared to this scenario. Concerning the above discussion and as seen in Fig (2), all the shapes of the free energy landscape (with a little tolerance for displacement) will practically have the same shape structure in the canonical view. Accordingly, the choice and use of each do not seem to have an advantage over the others. For this reason, we will henceforth utilize them under the designation of the free energy landscape, without loss of generality.
@@ 3.0.2 Kramer’s escape rate
An important point to focus on initially is the conceptual difference between this section and the previous one. In this section, we endeavor to interpret and study the dynamic sequence of a transfer process from a small black hole to a large black hole during a first-order phase transition, utilizing Kramer’s escape rate. This rate represents the escape velocity of particles under thermal fluctuations from a known local minimum in the potential.
We will also examine whether conditions could arise that reverse the transfer process from a small to a large black hole. Typically, we expect a transition from an unstable local minimum to a stable global minimum. However, the question arises whether there is a range in which, due to environmental conditions such as temperature or pressure changes, the likelihood of a reverse process, namely a transfer from a large black hole to a small black hole, precedes the primary and natural process.
For this particular scenario, we focus our attention on the frame-by-frame phase transition sequence based on the escape rate. An important point to note is that we are studying the dynamics, not the statics, of the frames. These frames are temporally continuous, meaning each frame at a given moment possesses its own physical concepts. However, these concepts must be interpretable in a way that they align with the overall understanding of the preceding and succeeding frames, ultimately corroborating the general perspective.
In Fig(5), we approach a state at the onset of phase transition where the chaotic process of phase transition has formed due to primary reasons, meaning the $\gamma$ minimum has found its initial form (as shown in Fig(5b)), although, it has not yet become a global minimum. The small black hole (in the local minimum potential $\alpha$) has begun its intense movement towards the large black hole, considering the formation of a secondary minimum and prevailing conditions, which is clearly observed in Fig(5c). The probability of transition $\alpha\rightarrow\gamma$ is significantly higher than the other transition, Fig(5c), which is why the dominant transfer seems to be entirely from $\alpha\rightarrow\gamma$. Since we are at the beginning of this transition, it seems that the big black hole has not yet formed, or in other words, it lacks enough particles for the formation of reverse processes. Therefore, as it is evident from the Fig(5d), the structure is under the complete control of the transition from a small black hole to a large one, and $\Delta r_{k}$ is completely positive.
Figure 5: t versus x and the First order Phase Transition region (FPT) in Fig (5a),  (5b): The graph of $u(r)$ against x for free parameters and the FPT points,   (5c): comparing $r_{k}$ with respect to $x$ for($\alpha\rightarrow\gamma$) and ($\gamma\rightarrow\alpha$) and the coordinate of contact point   Fig (5d): $\Delta r_{k}$ ($r_{k}(\alpha\rightarrow\gamma)-r_{k}(\gamma\rightarrow\alpha)$) with respect to $x$
In the subsequent stage, we examine a state midway through the phase transition, where the minimum $\gamma$ has assumed a relatively complete form (as depicted in Fig(6a)), and although it has not yet become a global minimum, it is on par in depth with the alpha minimum. If we were to consider this statically and in a single-frame format, it would appear that the escape rate and the probability of movement from the local beta maximum in both directions (i.e., towards the small or the large black hole) should be equal. However, it is important to note that on one hand, the environmental conditions imposing the phase transition on the system, and on the other hand, the shrinking radius —which is somewhat contrary to normal physical laws— still, make the probability of transition $\alpha\rightarrow\gamma$ greater than the other transition, as shown in Fig(6b). For this reason, the dominant transfer seems to continue to be from $\alpha\rightarrow\gamma$. Nevertheless, as is evident in Fig (6b), a collision point has formed between the escape rate diagrams, and beyond this contact point, it is clear that the reverse process, albeit in a very small region at the end of the diagram, predominates over the direct process. It appears that as we approach the end of the phase transition, a negative region (Fig (6c,6d)) is forming which somehow, by reinforcing feedback processes, sets the stage for a change in the initial conditions and ultimately halts the phase transition.
Figure 6: (6a): The $u(r)$ against x for free parameters and the FPT points,   (6b): comparing $r_{k}$ with respect to $x$ for($\alpha\rightarrow\gamma$) and ($\gamma\rightarrow\alpha$) and the coordinate of contact point   (6c):$\Delta r_{k}$ with respect to $x$   (6d): The enlarged detail of the right end of the Fig (6c)
Ultimately, we turn our attention to a state near the end of the phase transition pathway, where the $\gamma$ has taken on a relatively complete shape (as illustrated in Fig(7a)) and has become a global minimum. Although the environmental conditions that are imposing the phase transition on the system continue to make the probability of transition $\alpha\rightarrow\beta$ greater than the other transitions, as depicted in Fig(7b). However, as can be seen from Fig (7b), the coordinates of the contact point have moved significantly upwards, to the extent that the probability of the reverse process cannot be ignored in comparison to the direct process, and now a larger region is dominated by this process, as shown in Fig(7c). As we observed throughout this sequence, the closer we got to the end of the phase transition, the contact point advanced towards, increasing the probability of the reverse transition. It seems that this reverse process may, in some way, be the system’s reaction to prevent the phase transition process from becoming uncontrollable and the radial growth that might overshadow the stability of the black hole. It may also be hypothesized that the black hole, to prevent the phase transition process from becoming uncontrollable, instead of a direct and sudden transition from the small black hole to the large black hole (akin to a completely damped oscillation), performs this process in the form of a slowly damped, quasi-oscillatory manner.
@@
Figure 7: (7a): The $u(r)$ against x for free parameters and the FPT points,   (7b): comparing $r_{k}$ with respect to $x$ for($\alpha\rightarrow\gamma$) and ($\gamma\rightarrow\alpha$) and the coordinate of contact point   (7c):$\Delta r_{k}$ with respect to $x$
@@
where q is the charge of the black hole, ”a” is related to the density of clouds of strings and comes from the energy-momentum tensor of the clouds of strings and $g$ 666 Gauss-Bonnet coupling constant is displayed with g instead of $\alpha$ to avoid mistakes with clouds of strings parameter is a positive Gauss-Bonnet coupling constant.
@@ 4.0.2 Kramer’s escape rate
In order to examine the escape rate in this black hole, given that the motion trajectory is analogous to the previous state, we will delineate only two scenarios to prevent redundancy: one at the onset of phase transition and the other in proximity to the conclusion of phase transition. We endeavor to eschew similar descriptions as much as possible.
Figure 8: t versus x and the First order Phase Transition region (FPT) in Fig (8a),  (8b): The graph of $u(r)$ against x for free parameters and the FPT points,   (8c): comparing $r_{k}$ with respect to $x$ for($\alpha\rightarrow\gamma$) and ($\gamma\rightarrow\alpha$) and the coordinate of contact point   Fig (8d): $\Delta r_{k}$ with respect to $x$
As evident from Fig(8d), just like before, at the initial of phase transition of the process is predominantly characterized by a direct transition. As observed in the diagram, the positive region is effectively the dominant area and no reactive structure appears until the end of the process.
Figure 9: (9a): The $u(r)$ against x for free parameters and the FPT points,   (9b): comparing $r_{k}$ with respect to $x$ for($\alpha\rightarrow\gamma$) and ($\gamma\rightarrow\alpha$) and the coordinate of contact point   (9c):$\Delta r_{k}$ with respect to $x$
As illustrated in Fig(9), with the progression of the phase transition process and as we approach the end, although the primary contribution is still from a direct transition, as depicted in Fig(9d), but the negative region has undergone significant growth and has become comparable to the positive region. It appears that, similar to the previous case, the reactive processes have become more potent in near the end of the process.
@@ 5 Discussion and comparison
As observed in Figs (5b),(6a),(8b), (7a) and (9a), our black hole initially resides in a stable local minimum,$\alpha$, which globally is considered an unstable minimum. Consequently, as one would naturally expect, with the slightest thermal perturbation, and corresponding pressure, the system increasingly tends to transition towards the global minimum,$\gamma$, signifying an increase in radius.
In both the NLM-C-Q-PFD and EGB-YM-CS models, upon examining the escape rate diagrams, we observe that once the radius exceeds a certain threshold, a small domain emerges where the reverse process forms and prevails. In other words, with the dominance of $r_{k}(\gamma\rightarrow\alpha)$ over $r_{k}(\alpha\rightarrow\gamma)$, a condition arises allowing black holes to transition from $\gamma$ to $\alpha$. Notably, over time and the sequence of frames, this seemingly negligible and very small region at the onset of phase transition gradually transforms into a significant area in the final frames. At first glance, the transition from $\gamma$ to $\alpha$ may seem less compatible with physical conditions because, within the same energy structure and without topological changes in the energy function (i.e., maintaining the gamma state as the global minimum and alpha as the local minimum), the system prefers to revert to the ( $\alpha$ ) state with a smaller radius, leading to a decrease in entropy and an increase in energy level. This situation can be termed ’improbable’ or ’less likely’ from a physical standpoint. However, on the other hand, from a Kramers’ perspective on the reverse process, it can be stated that particles situated in a potential well under thermal fluctuations can always escape with a specific probability coefficient (which is a function of the potential depth). Therefore, from the perspective of particles in the gamma potential, this movement is not considered an improbable phenomenon. Moreover, if we consider the transition from a smaller to a larger black hole as a general action, a natural reaction should form against it. The increase in radius and the consequent decrease in energy level is a recognized form of this reaction, leading to the reality that the escape rate, after reaching a peak, undergoes a decreasing trend until the probability becomes zero. But does this zero probability occur at any desired radius?
These ideas, coupled with the observation that as we approach the terminal point of phase transition, we witness an increased probability of the reverse process, have led us to propose that perhaps this reverse trend is, in fact, a natural response of the black hole to phase transition. It may indeed be an effort to limit and control the phase transition to prevent excessive radial growth and uncontrolled instability, ensuring that the process of moving towards a larger black hole concludes under conditions that maintain the stability of the final structure and do not disrupt the viability of the black hole.